# DOMBA: Double Model Balancing for Access-Controlled Language Models via Minimum-Bounded Aggregation

### v1

@@ Results
The results with two epochs of training are presented in Table 1. In terms of utility, FT-ALL performed the best across both metrics, as expected. Among the secure and access-controlled models, D-I-H achieved the highest utility on the HOPPL metric, while DOMBA achieved the highest utility on the AUPPL metric, with a substantial gap compared to secure models, highlighting the importance of the DOMBA-FT step. Comparing access-controlled models, DOMBA exhibited substantially better utility than PER-AL across both metrics and datasets.
@@
Figure 2 shows the worst-case and 99th percentile exposure of models employing different aggregation methods on the recipe and movie review datasets for 1, 2, and 4 training epochs. The maximum exposure of DOMBA, D-I-H, and D-I-M is 4.69. In comparison, SUBMIX reaches a maximum exposure of 8.5e4 and AGG-A reaches a maximum exposure of 1.3e10. We observe that DOMBA’s 99th percentile exposure is similar to its maximum exposure, supporting the theoretical bound established by our analysis (Theorem 2). Regarding the effect of the nubmer of epochs, increasing it generally leads to higher exposure (except for SUBMIX’s 99th percentile exposure on the recipe dataset). However, the increase in exposure for DOMBA, D-I-H, and D-I-M is moderate compared to AGG-A for both datasets, while for SUMBIX, the change in exposure is inconsistent between the two datasets.
@@ Discussion
Our results show that D-I-H and D-I-M achieve a good utility-security trade-off. This suggests that our method may be suitable for general privacy-preserving purposes beyond access control. In future research, it will be interesting to develop a variation of DOMBA for non-access-controlled private datasets and compare its performance to that of state-of-the-art privacy-preserving methods which are not focused on the access-control scenario.
One limitation of DOMBA is that its security cannot be increased further (assuming minimum is used as the aggregation function). This is opposed to DP methods which can achieve any security level (with a cost on utility) by adjusting a privacy parameter. Future research could explore hybrid approaches, combining DOMBA with DP techniques to offer security beyond DOMBA’s current maximum level of security.
Additionally, DOMBA relies on a strict separation of access levels into two distinct partitions without shared sensitive information. Such separation could be challenging to implement in some scenarios. To make DOMBA more robust to sensitive information shared between access levels, future research could explore the separation of the access levels into more than two partitions.
Resource overhead is incurred with DOMBA’s deployment due to the use of two LLMs instead of one, which may be impractical for some applications. One potential solution is to employ DOMBA as a teacher model to train a student model via knowledge distillation (Xu et al. 2024), where the student model serves as a deployed model mimicking DOMBA.
@@ Conclusion
In this paper, we proposed DOMBA, a novel approach for training and deploying access-controlled LLMs with high utility. We formalized the concept of exposed secrets and bounded DOMBA’s exposure. We evaluated DOMBA’s performance on two access-controlled datasets, reflecting real world organizations’ needs. Our evaluation showed that DOMBA achieves a better security-utility trade-off than existing methods, across both datasets, two utility metrics and four security metrics. Finally, we believe that the principles of min-bounded aggregation and relative probabilities, which serve as DOMBA’s core, have substantial potential to serve as foundational elements in a wide range of future machine learning research, extending beyond the scope of security.

### v2

@@ Results
The results with two epochs of training are shown in Table 1. As expected, FT-ALL achieved the best utility across both metrics. Among secure and access-controlled models, D-I-H had the highest HOPPL utility, while DOMBA excelled on the AUPPL metric, demonstrating the value of the DOMBA-FT step. Comparing access-controlled models, DOMBA substantially outperformed PER-AL across both datasets and both utility metrics.
@@
Regarding the impact of the choice of min-bounded function, D-I-M achieved better security and worse utility compared to D-I-H, as expected. We also tested replacing the min-bounded function used in DOMBA from minimum to harmonic mean and the effect was similar.
@@
Figure 2 shows the worst-case and 99th percentile exposure of models employing different aggregation methods on the recipe and movie reviews datasets for 1, 2, and 4 training epochs. The maximum exposure of DOMBA, D-I-H, and D-I-M is 4.69. In comparison, SUBMIX reaches a maximum exposure of 8.5e4 and AGG-A reaches a maximum exposure of 1.3e10. We observe that DOMBA’s 99th percentile exposure is similar to its maximum exposure, supporting the theoretical bound established by our analysis (Theorem 2). Regarding the effect of the number of epochs, increasing it generally leads to higher exposure. However, the increase in exposure for DOMBA, D-I-H, and D-I-M is moderate compared to AGG-A for both datasets, while for SUMBIX, the change in exposure is inconsistent between the two datasets.
@@ Discussion
To analyze DOMBA, we developed the token exposure framework, offering an alternative to DP for theoretically analyzing the security of language models. The key idea behind token exposure is to evaluate the difference between a safe model and a potentially unsafe one. In contrast, DP evaluates the differences between models trained on nearly identical datasets (in both approaches, a smaller difference indicates stronger security). Additionally, token exposure utilizes relative probabilities over absolute ones. We find the token exposure approach more intuitive and practical for analyzing DOMBA. Future work could further explore the token exposure framework and investigate its relationships and potential synergies with DP.
DOMBA relies on a strict separation of access levels into two distinct partitions that do not share sensitive information. Achieving this separation requires the graph of shared secrets between access levels to be disconnected. We hypothesize that this condition is usually met, as organizations often handle documents across diverse topics. Future research could validate this hypothesis and investigate approaches for scenarios where the graph remains connected.
While the training time of DOMBA scales linearly with the dataset size, similar to training a single LLM, and storage requirements can be kept low by using LORA, inference incurs additional resource overhead due to the deployment of two LLMs instead of one. This overhead may render the approach impractical for certain applications. One potential solution is to employ DOMBA as a teacher model to train a student model via knowledge distillation (Xu et al. 2024; Papernot et al. 2017), where the student model serves as a deployed model mimicking DOMBA.
@@ Conclusion
In this paper we proposed DOMBA, a novel approach for training and deploying access-controlled LLMs with high utility. We formalized the concept of exposed secrets by developing the token exposure framework and bounded DOMBA’s exposure. We evaluated DOMBA’s performance on two access-controlled datasets, mimicking real world organizations’ needs. Our evaluation showed that DOMBA achieves a better security-utility trade-off than existing methods, across both datasets, two utility metrics and four security metrics. Finally, we believe that the principles of min-bounded aggregation and relative probabilities, which serve as DOMBA’s core, have substantial potential to serve as foundational elements in a wide range of future machine learning research, extending beyond the scope of security.
Acknowledgments
The authors would like to thank Dr. Avi Segal for their valuable insights and thoughtful feedback, which contributed to the refinement of this manuscript.